OpenStudy (anonymous):

A researcher wants to conduct a genetic study using 15 randomly-selected volunteers. He has a volunteer pool of 30, composed of 15 males and 15 females. What is the probability that the randomly-selected group will be all male OR all female? A. 1/155,117,520 B. 1/77,558,760 C. 1/310,235,040 D. 1/450

OpenStudy (anonymous):

@AndrewKaiser333

OpenStudy (andrewkaiser333):

I honestly do not know much for probability never had to go through it yet this year. my guess would be D. or B. but i would rather you ask hero, thomaster or sattelite so since i said that i will recommend them now @thomaster @Hero @satellite73

OpenStudy (anonymous):

@thomaster @Hero @satellite73

OpenStudy (andrewkaiser333):

sorry I am kinda stumped because I never went through it like I said but i gave you my guesses and recommended 3 great math people who have been through this subject I think.

OpenStudy (anonymous):

thank you

OpenStudy (andrewkaiser333):

np hun

OpenStudy (anonymous):

@ajaylounsber

OpenStudy (anonymous):

@joanip_

hero (hero):

I completed the calculation

OpenStudy (anonymous):

I don't know, sorry :s

OpenStudy (anonymous):

@Hero okay what did u come up with

hero (hero):

Basically: The probability that the 1st selected will be male is \(\dfrac{15}{30}\) The probability that the 2nd selected will be male is \(\dfrac{14}{29}\) The probability that the 3rd selected will be male is \(\dfrac{13}{28}\) And so on The probability that the 15th selected will be male is \(\dfrac{1}{16}\) So the probability that the first 15 selected will be all male is \(\dfrac{15}{30} \dot\ \dfrac{14}{29} \dot\ \dfrac{13}{28}...\dfrac{1}{16} = \dfrac{1}{155117520}\) The probability that the first 15 selected will be all female is \(\dfrac{15}{30} \dot\ \dfrac{14}{29} \dot\ \dfrac{13}{28}...\dfrac{1}{16} = \dfrac{1}{155117520}\) Therefore the probability that the first 15 selected will be all male or all female is \( \dfrac{1}{155117520} + \dfrac{1}{155117520}\)

OpenStudy (anonymous):

thank you soooo much ! do u feel like answering another question? Its answered but i want to make sure its right @Hero